OpenStudy (anonymous):

Which equation below is not an equation that represents the graph shown below? line through the points 0 comma 1 and 1 comma negative 2 Choose one answer. a. y + 2 = -3(x - 1) b. 3x + y =1 c. y - 1 = -3(x + 2)

OpenStudy (anonymous):

d. y = -3x + 1

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (0,1) x1 = 0 and y1 = 1 (1,-2) x2 = 1 and y2 = -2 sub slope(m) = (-2-1) / (1 - 0) slope (m) = -3/1 which is -3 y = mx + b slope(m) = -3 (1.-2) x = 1 and y = -2 -2 = -3(1) + b -2 = -3 + b -2 + 3 = b 1 = b y = -3x + 1(correct) and 3x + y = 1 (correct) y - y1 = m(x - x1) (1,-2) x1 = 1 and y1 = -2 slope = -3 sub y - (-2) = -3(x - 1) y + 2 = -3(x - 1) (correct) your answer is C

OpenStudy (anonymous):

Thank you ! Can you help me with this one i have a hard time understing it :[[ Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = one thirdx - 3. Choose one answer. a. y = 3x - 14 b. y = one thirdx + 2 c. y = -3x + 22 d. y = -one thirdx + 8

OpenStudy (texaschic101):

y = 1/3x - 3 slope is 1/3. A parallel line will have the same slope. y - y1 = m(x - x1) y - 4 = 1/3(x - 6) y - 4 = 1/3x - 1/2 y = 1/3x - 1/2 + 4 y = 1/3x - 1/2 + 8/2 y = 1/3x + 7/2 ===> answer my answer does not match your answer choices However, the only answer choice that comes close is B, because it has the same slope.